UT Feb 03 2021

Night Center: 00:33 - handover from PEPSI 23:45

Twilight Morning 05:49

Night Plan (second half)

  • Reconfigure for LUCI-Bino at 23:45 - authorize LUCI-1
  • AO Activities
    • OSCO after reconfiguring - 0.5hr - to be done as long as loop can be closed
    • 12:30 AM (local time) 2020 SO LUCI-1 AO - 1hour
      • If < 1.2" and thin clouds proceed to 2020 SO LUCI-1 Full AO
      • Ace Track script, need to collimate ARGOS-style beforehand
    • AO Bin 1,3 checks - AO star at dfferent rotator angles - 1 hour
    • ESM Bin 2 checks - observe M67 or M92 H-band at different rotator angles - 1 hour
  • If AO cannot be done - reconfigure for LBCs - either poor weather or unable to move SX M3
    • 2020 SO with LBCs (backup to program above) - 1hour - target is brightest after midnight
  • If weather is (<1.2" and thin clouds or better) and cannot move to LUCI (and have already done 2020 SO with LBCs) - reconfigure for MODS
    • MODS science programs
      • ULIRG velocity dispersions (Rothberg) - 1.5hr per target
      • IMBH Candidate (Rothberg) - 1.5hr
        • target visible after 3:30am
  • Else If poor weather and cannot move M3 after 2020 SO with LBCs - reconfigure for MODS
    • MODS backup science program - Local Mergers Velocity dispersions (Rothberg) - 1.5 hr per target
    • MODS Active Optics parameter changes (J. Hill) - (1 hr in parallel)
  • Else If poor weather but can move M3 to LUCI - after 2020 SO with LBCs - reconfigure for LUCIs
    • LUCI backup science program - ICRF3 (Rothberg) 0.5 hr per target
    • LUCI Active Optics parameter changes (J. Hill) - (1 hr in parallel)
Script Location: /home/lbto/Service_Observing/D_E_Feb_Mar_2021

Discussion with I. Illyn - the transit time for tonights target is actually 5am loacl time - PI may have made an error again on the transit times and observing request.

Night Log

Pre-Sky Prep

UT 23:12 (Feb 1) - Conditions look pretty poor - will bring up facility instruments shortly.

UT 01:43 - bringing up LBCs on ROBS64 - will use OBS2 for LUCI and MODS

UT 01:49 - taking test biases with LBCs

-look about right

UT 01:53 - waking up MODS - weren't connceted to TCS - reconnecting now

UT 01:58 - MODS awake, taking sievesnape -- MODS-2 seems stuck on CALMODE

-CTRL C'd and ran again for MODs2

-looks OK

UT 02:05 - bringing up LUCI GUIs

-init all , test dakrs 2.51sec

L1 0001-0002

L2 0001-0002

-looks OK

UT 02:08 - three facility instruments up and reay

UT 02:51 - can barely see the Moon through the clouds

UT 05:37 - OSA reports clear enough to open - need to contact I.Ilyn - looks like going to open


ON-Sky

UT 06:03 - Open but clouds coming back and having some issues with mirrors

UT 06:10 - Closing up again - I.Ilyn will hand over 20min earlier

-M3 SX successfully moved from PEPSI to LUCI!

UT 08:07 - still closed - thick clouds overhead

UT 09:41 - still closed, thick clouds overhead - short periods of holes in clouds from time to time

UT 12:25 - powering off LBCs and logging out of ROBS64

UT 12:26 - putting MODS to sleep


Closed Dome Cals

MODS

UT 06:42 - taking MODS cals

Arcs Dual Grating

M1B 0002-0004

M1R 0002-0004

M2B 0002-0004

M2R 0002-0004

- MODS1 blue last etxposure lamps timed out - hit Retry

-nneded to run red fitsflush on MODS2B

Slitless Pix Flats Dual Grating

M1B 0005-0014

M1R 0005-0009

M2B 0005-0014

M2R 0005-0009

8K biases

M1B 0015-0019

M1R 0010-0014

M2B 0015-0019

M2R 0010-0014

3K Biasses

M1B 0020-0024

M1R 0015-0019

M2B 0020-0024

M2R 0015-0019

LUCI

-N30 AO Imaging flats JHKKs

UT 12:51 - starting cals...and we have a camera wheel error on LUCI-1

-moving from N375 to N30 at park and zenith

wheel is at -157 now

-init camera , trying to find neg limit, moved to -857

-using IMGUI to move to N30

-moved to -228

-strange....simply copying the J-band template from the library created a script that has J,H,Ks,K and FeII.. putting a pause after K-band

N30 field stop is way off - aborted lights on J-band

L1 0003-0007 - do not use

L1 0008-0011 N30 alignment

UT 13:07 - restarting flat script

L1 0012-0021 - J-band

L1 0022-0031 - H-band

L1 0032-0041 - Ks-band

L1 0042-0051 - K-band

UT 13:31 - putting LUCI-1 into safe config (LUCI-2 already in safe config)

DONE

-- BarryRothberg - 01 Feb 2021
Topic revision: r10 - 03 Feb 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback